Suppose we experiment with an Atwood Machine consisting of two masses m


and 2m connected by an inextensible non-slip string running over a pulley of


radius a and mass 4m, as shown below.


(a) Determine the total kinetic energy of the system. (6)


(b) Find the gravitational potential energy of the system.

1
Expert's answer
2022-03-22T11:00:31-0400

"a=g\frac{m_2-m_1}{m_2+m_1}=3.27~\frac m{s^2},"

"I=2mr^2,"

"\omega=\frac ar,"

"E=\frac{I\omega^2}2=10.7mr^2~J."